High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costingReg 1 Req 2 Req 3A Req 3B Compute the Predetermined Overhead Rates used in the Molding Department and the Painting DepartmentReq 1 Req 2 Req ЗА Req 3B Compute the total overhead cost applied to Job 205. (Round Predetermined overhead rate to 2 decimReq 1 Reg 2 Req ЗА Req 3B What would be the total manufacturing cost recorded for Job 205? (Round Predetermined overhead ratComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req ЗА Req 3B If the job contained 35 units, w
Answer #1

1) Molding department = (210000/84000+2.40) = 4.90 per MH

Painting department = (479320/52100+4.40) = 13.60 per DLH

2) Total overhead cost = (400*4.90+125*13.60) = 3660

3) Total manufacturing cost

Molding Painting Total
Direct material 934 1140 2074
Direct labor 710 970 1680
Overhead 1960 1700 3660
Total 3604 3810 7414

3b) Unit product cost = 7414/35 = 211.83